A ray of light in air-incident at an angle of 45° on the boundary with water. Calculate the angle of refraction if the relative refractive index of air is 0.55?

1
Expert's answer
2022-02-08T10:12:19-0500

Given:

"\theta_i=45^\circ"

"n=0.55"

The Snell-Decart law says

"\frac{\sin\theta_i}{\sin\theta_r}=\frac{1}{n}"

The angle of refraction

"\theta_r=\sin^{-1}(n\sin\theta_i)"

"\theta_r=\sin^{-1}(0.55*\sqrt{2}/2)=23^\circ"